OpenStudy (anonymous):

3W = 4 and 1/5ths (fraction not division) solve for the variable

OpenStudy (anonymous):

turn 4 and 1/5 into just one fraction. 4x5+1 =21 so it will be 21/5 \[3W=\frac{21}{5}\] \[W=\frac{21}{5}\div3\] what does W equal to now?

OpenStudy (anonymous):

seven fifths thanks

OpenStudy (anonymous):

yeh, medal thanks lol.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

:)